# Biblio.bib

### v2

@@ 2.4. Weitzenböck formula
The last tool we will use is a Weitzenböck formula. Consider the symmetric $2$-tensor $S$ as $1$-form with values in $T^{*}M$. Denote by $d^{\nabla}$ the differential induced by the Levi Civita connection. Using [Bes, (12.69)] gives for any $S\in C^{\infty}(M;S^{2}T^{*}M)$
@@
[…]
@@
$\langle R^{\circ}(S_{0}),S_{0}\rangle=\Lambda_{2}^{n}\langle\pi_{2}^{*}R^{\circ}(S_{0}),\pi_{2}^{*}S_{0}\rangle_{L^{2}(SM)}$ =$=\begin{cases}\|S_{0}\|^{2}\\ \|S_{0}\|^{2}-3\langle S_{0}\circ J,S_{0}\rangle\\ \|S_{0}\|^{2}-3\sum_{i=1}^{3}\langle S_{0}\circ J_{i},S_{0}\rangle\\ \langle R^{\circ}(S_{0}),S_{0}\rangle.\end{cases}$
@@
[…]
@@ 3.1. Re-writing the Hessian
The first step consists in using the fact that $g_{0}$ is a locally symmetric metric to obtain an explicit formula for $d^{2}\Phi(g_{0})$. The computation of the Hessian is done by Flaminio in [Fla, Proposition 5.1.1] for real hyperbolic metrics and we adapt it to the other classes of locally symmetric spaces. We prove the following theorem.
@@
(3.1) $\langle d^{2}\Phi(g_{0})S,S\rangle_{L^{2}(S^{2}T^{*}M)}=\langle\Pi_{m(g_{0})}^{g_{0}}Q(S),Q(S)\rangle-\langle\pi_{m(g_{0})}^{*}Q(S),1\rangle_{L^{2}(SM)}^{2}$
where $m(g_{0})\in\mathbb{N}$ and $Q:C^{\infty}(M;S^{2}T^{*}M)\cap\mathrm{Ker}(D_{g_{0}}^{*})\to C^{\infty}(M;S^{m(g_{0})}T^{*}M)$ is a differential operator. More precisely, with the notations introduced in Section 2,
  • If $g_{0}$ is real hyperbolic, then $m(g_{0})=2$ and
    (3.2) $Q(S)=-\frac{1}{2}S+\frac{1}{4}\nabla^{*}\nabla S+\frac{1}{2}(\mathrm{tr}_{g_{0}}S)g_{0}.$
  • If $g_{0}$ is complex hyperbolic then $m(g_{0})=4$ and
@@
  • If $g_{0}$ is quaternionic hyperbolic, then $m(g_{0})=4$ and  [^3] Again, it is a statement of the theorem that the formula below is defined globally and not just locally.
@@
  • If $g_{0}$ is octonionic hyperbolic, then $m(g_{0})=4$ and
    (3.5) $Q(S)=L\big{(}\frac{1}{3}\operatorname{tr}(S)g_{0}+\frac{1}{6}S-\frac{1}{6}R^{\circ}(S)+\frac{7}{24}\nabla^{*}\nabla S\big{)}-\frac{1}{24}\mathrm{Sym}(R^{\circ}(\nabla^{2}S)).$
@@
Consider a smooth path of metrics $(g_{\varepsilon})$ and write $S=\partial_{\varepsilon}g_{\varepsilon}|_{\varepsilon=0}$. First, we can identify all sphere bundles $S_{g}M$ by the pullback by $\Psi_{g}:SM\to S_{g}M$ defined by $\Psi_{g}(x,v):=(x,v/|v|_{g})$. Next, for $\varepsilon$ small enough, the Anosov stability theorem ensures the existence of a homeomorphism $\Psi_{\varepsilon}\in C^{\alpha}(SM,SM)$ which gives an orbit equivalence:
(3.6) $\Psi_{\varepsilon}(\varphi_{t}(z))=\varphi^{\varepsilon}_{\kappa_{\varepsilon}(z,t)}(\Psi_{\varepsilon}(z))$
where the index $\varepsilon$ indicates that we consider the corresponding objects for the metric $g_{\varepsilon}$ and $\kappa_{\varepsilon}$ is a time rescaling. Following Flaminio, we define
$\bar{\gamma}:(-\varepsilon_{0},\varepsilon_{0})\times\mathbb{R}\mapsto\bar{\gamma}(\varepsilon,t):=\pi(\Psi_{\varepsilon}(\varphi_{t}v)),$
where $\pi:SM\to M$ is the projection. This means that $\bar{\gamma}(\varepsilon,\cdot)$ is a $g_{\varepsilon}$ (non normalized) geodesic. We define $Y=\tfrac{d\bar{\gamma}}{dt}$ and $\xi=\tfrac{d\bar{\gamma}}{d\varepsilon}$. Define $U(t,\varepsilon):=U_{g_{\varepsilon}}(\Psi_{\varepsilon}(\varphi_{t}v))$ where $U_{g_{\varepsilon}}(v)$ is the second fundamental form of the stable horosphere for the metric $g_{\varepsilon}$ at $\pi_{\varepsilon}(v)\in M$, seen a $(1,1)$-tensor. Note that, as already noticed in [Fla, Page 582] and with the notations of (2.21), along a geodesic, one has $UY_{j}=-\lambda_{j}Y_{j}$. Our starting point is [Fla, Proposition A, $(C_{3})$, Lemma 3.3.11, Lemma 4.2.1] which gives [^4] We use Pesin’s entropy formula to simplify the topological entropy appearing in [Fla, Proposition A, $(C_{3})$] with the sum of negative Lyapunov exponents appearing in [Fla, Lemma 3.3.11]:
(3.7) $\frac{d^{2}}{d\varepsilon^{2}}\Phi(g_{\varepsilon})|_{\varepsilon=0}=\mathrm{Var}_{\mu_{\mathrm{Liou}}}^{g_{0}}(\operatorname{tr}(\nabla_{\xi}U)).$
Now, the goal is to write more explicitly the Hessian as a function of the infinitesimal variation $S=\partial_{\varepsilon}|_{\varepsilon=0}g_{\varepsilon}$. We now consider parallel fields $Y_{1}=Y,Y_{2},\ldots,Y_{n}$ as in (2.21). The Riccati equation satisfied by $U_{g}$ translates into a differential equation satisfied on the diagonal entries $B_{j}:=\langle(\nabla_{\xi}U)Y_{j},Y_{j}\rangle$. This is the content of [Fla, Lemma 4.3.1] and we have [^5] We are using a different convention for the sign of $R$, hence the sign difference with [Fla, Lemma 4.3.1]. :
(3.8) $\mathcal{L}_{Y}B_{j}-2\lambda_{j}B_{j}=-S(Y,Y)\lambda^{2}_{j}-g_{0}\left(\frac{\partial R_{\varepsilon}}{\partial\varepsilon}|_{\varepsilon=0}(Y,Y_{j})Y,Y_{j}\right).$
@@
(3.9) $\sum_{j=1}^{n}B_{j}=\operatorname{tr}(\nabla_{\xi}U)\approx\frac{1}{2}\left(\sum_{j=2}^{n}\lambda_{j}\right)S(Y,Y)+\frac{1}{2}\sum_{j=1}^{n}\mu_{j}g_{0}\left(\frac{\partial R_{\varepsilon}}{\partial\varepsilon}|_{\varepsilon=0}(Y,Y_{j})Y,Y_{j}\right),$
where $\mu_{1}=1$ and $\mu_{j}=\lambda_{j}^{-1}$ for $j>1$. Here, $\approx$ denotes the fact that the two functions are cohomologous, i.e their difference is of the form $Xu$ where $X$ is the geodesic vector field associated to $g_{0}$ and $u\in C^{\infty}(SM)$ and we noted that the terms $\mathcal{L}_{Y}B_{j}$ are co-boundaries. We note that (3.8) for $j=1$ gives that $g_{0}\left(\frac{\partial R_{\varepsilon}}{\partial\varepsilon}|_{\varepsilon=0}(Y,Y)Y,Y\right)\approx 0$ (recall from (2.21) that $\lambda_{1}=0$) and we can thus add it in the right hand side. Moreover, we can compute the diagonal entry $B_{1}$:
@@
To understand the second term in the right hand side of (3.9), we need to compute the variation of the different curvature tensors with respect to the metric. The computations are done in [Bes, Theorem 1.174] and more precisely, the variation of the $(3,1)$ curvature tensor $R_{g}$ defined in (2.19) is given by:
$\frac{\partial R_{\varepsilon}}{\partial\varepsilon}|_{\varepsilon=0}S(X,Y)Z=(\nabla_{Y}{\nabla^{\prime}_{g}}S)(X,Z)-(\nabla_{X}{\nabla^{\prime}_{g}}S)(Y,Z)$
@@
$g_{0}({\nabla^{\prime}_{g}}S(X,Y),Z):=g_{0}\left(\frac{\nabla^{g_{\varepsilon}}_{X}Y}{\partial\varepsilon}|_{\varepsilon=0},Z\right)=\frac{1}{2}(\nabla_{X}S(Y,Z)+\nabla_{Y}S(X,Z)-\nabla_{Z}S(X,Y)),$
where $S=\partial_{\varepsilon}|_{\varepsilon=0}g_{\varepsilon}.$ For sake of notations, we will denote by $C_{g_{0}}$ the $(3,0)$ tensor field
@@
$\sum_{j=2}^{n}\mu_{j}g_{0}\left(\frac{\partial R_{\varepsilon}}{\partial\varepsilon}|_{\varepsilon=0}(Y,Y_{j})Y,Y_{j}\right)=\sum_{j=2}^{n}\mu_{j}\left(\nabla_{Y_{j}}(C_{g_{0}}S)(Y,Y,Y_{j})-\nabla_{Y}(C_{g_{0}}S)(Y_{j},Y,Y_{j})\right).$
@@
$\sum_{j=1}^{n}\mu_{j}g_{0}\left(\frac{\partial R_{\varepsilon}}{\partial\varepsilon}|_{\varepsilon=0}(Y,Y_{j})Y,Y_{j}\right)\approx\sum_{j=1}^{n}\mu_{j}\nabla_{Y_{j}}(C_{g_{0}}S)(Y,Y,Y_{j}).$
@@
$\nabla_{Y_{j}}(C_{g_{0}}S)(Y,Y,Y_{j})=\nabla_{Y_{j},Y}^{2}S(Y,Y_{j})-\frac{1}{2}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y).$
We use the Ricci identity and (2.21), to re-write the first term:
@@
$=\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})+\lambda_{j}^{2}S(Y_{j},Y_{j})+S(R(Y_{j},Y)Y_{j},Y).$
This means that
$\nabla_{Y_{j},Y}^{2}S(Y,Y_{j})=\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})+\lambda_{j}^{2}S(Y_{j},Y_{j})-\lambda_{j}^{2}S(Y,Y).$
@@
$-\frac{1}{4}\sum_{j=1}^{n}\mu_{j}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)=-\frac{1}{4}\sum_{j=1}^{n}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)=\frac{1}{4}\pi_{2}^{*}(\nabla^{*}\nabla S),$
@@
$-\frac{1}{2}\sum_{j=1}^{n}\mu_{j}\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})=-\frac{1}{2}\sum_{j=1}^{n}\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})=\frac{1}{2}\pi_{2}^{*}(D_{g_{0}}D_{g_{0}}^{*}S)\approx 0$
@@
Putting everything together gives $V(S)=\pi_{2}^{*}Q(S)$ where $Q$ is defined in (3.2). Now, using [GuKnLef, Equation (2.6)]:
$\mathrm{Cov}^{\varphi_{t}^{0}}_{\mu_{\mathrm{Liou}}^{g_{0}}}(\pi_{2}^{*}Q,\pi_{2}^{*}Q)=\langle\Pi_{2}^{g_{0}}Q,Q\rangle-\langle\pi_{2}^{*}Q,1\rangle_{L^{2}(SM)}^{2}$
and this gives (3.1). Note that $Q$ corresponds to the operator $T$ used by Flaminio in [Fla, Proof of Proposition 1.3.3].
@@
$-\frac{1}{4}\sum_{j=1}^{n}\mu_{j}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)$ $=-\frac{1}{4}\sum_{j=1}^{n}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)+\frac{1}{8}\nabla_{JY,JY}S(Y,Y)$
$=\frac{1}{4}\pi_{2}^{*}(\nabla^{*}\nabla S)+\frac{1}{8}\nabla_{JY,JY}S(Y,Y),$
@@
$\pi_{4}^{*}(\nabla_{JY,JY}S(Y,Y))=\pi_{4}^{*}\big{(}(\nabla\nabla(S\circ J))\circ J\big{)}=\pi_{4}^{*}\big{(}\mathrm{Sym}((\nabla\nabla(S\circ J))\circ J)\big{)}.$
@@
Here $\hat{X}_{i}$ denotes the $3$-uple where we left out $X_{i}$. In particular one has, using (2.10) and (2.25),
@@
Now, using the fact that $J$ is parallel, we see that $\nabla_{v,Jv}S(v,Jv)=Y(\nabla S_{x}(Jv,v,Jv))$ is a co-boundary and we can thus ignore this term when computing the Hessian. Finally, we use [GuKnLef, Equation (2.6)] and (3.10) to obtain (3.3).
@@
by (2.41). This also shows that is term is independent of the choice of the base $(J_{1},J_{2},J_{3})$ and thus globally defined. The last term $-\frac{1}{2}\sum_{j=1}^{n}\mu_{j}\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})$ is still a co-boundary by the exact same computation, using the fact that the $Y_{i}$ are parallel. We are left with understanding the middle term:
@@
$\pi_{4}^{*}\mathrm{Sym}\left(\sum_{j=1}^{3}\nabla_{J_{j+1}Y,J_{j+1}Y}S(Y,Y)\right)=-\pi_{4}^{*}\left(\sum_{i=1}^{3}D_{g_{0}}^{J_{i}}D_{g_{0}}^{J_{i}}S\right).$
@@
Quotient of $\mathbb{O}\mathbb{H}^{n}$. This time the first term gives
@@
where we used (2.52). The second term is still a coboundary and the last one is computed using using the fact that the $Y_{j}$ are parallel

### v3

@@ 2.5. Weitzenböck formula
The last tool we will use is a Weitzenböck formula. Consider the symmetric $2$-tensor $S$ as $1$-form with values in $T^{*}M$. Denote by $d^{\nabla}$ the differential induced by the Levi Civita connection. Using [26, (12.69)] gives for any $S\in C^{\infty}(M;S^{2}T^{*}M)$
@@
Except if stated otherwise, the notation $\langle\cdot,\cdot\rangle$ will denote the scalar product on $L^{2}(S_{0}^{2}T^{*}M)$ in the following proof. Suppose $S_{0}\in C^{\infty}(M;S^{2}T^{*}M)$ is trace-free, then equation (2.7) gives
@@
$\langle R^{\circ}(S_{0}),S_{0}\rangle=\Lambda_{2}^{n}\langle\pi_{2}^{*}R^{\circ}(S_{0}),\pi_{2}^{*}S_{0}\rangle_{L^{2}(SM)}$ $=\begin{cases}\|S_{0}\|^{2}\\ \|S_{0}\|^{2}-3\langle S_{0}\circ J,S_{0}\rangle\\ \|S_{0}\|^{2}-3\sum_{i=1}^{3}\langle S_{0}\circ J_{i},S_{0}\rangle\\ \langle R^{\circ}(S_{0}),S_{0}\rangle.\end{cases}$
@@
$0\leq\langle$ $\nabla^{*}\nabla S_{0},S_{0}\rangle-\langle R^{\circ}(S_{0}),S_{0}\rangle+\langle S_{0}\circ\mathrm{Ric},S_{0}\rangle.$
@@ 3.1. Re-writing the Hessian
The first step consists in using the fact that $g_{0}$ is a locally symmetric metric to obtain an explicit formula for $d^{2}\Phi(g_{0})$. The computation of the Hessian is done by Flaminio in [13, Proposition 5.1.1] for real hyperbolic metrics and we adapt it to the other classes of locally symmetric spaces. We prove the following theorem.
@@
(3.1) $\begin{split}\langle d^{2}\Phi(g_{0})S,S\rangle_{L^{2}(S^{2}T^{*}M)}&=\mathrm{Var}(\pi_{4}^{*}Q(S)).\end{split}$
where $Q:C^{\infty}(M;S^{2}T^{*}M)\cap\mathrm{Ker}(D_{g_{0}}^{*})\to C^{\infty}(M;S^{4}T^{*}M)$ is a differential operator. More precisely, with the notations introduced in Section 2,
  • If $g_{0}$ is real hyperbolic, then
    (3.2) $Q(S)=L\big(\frac{1}{4}\nabla^{*}\nabla S+\frac{1}{2}(\mathrm{tr}_{g_{0}}S)g_{0}-\frac{1}{2}S\big).$
  • If $g_{0}$ is complex hyperbolic, then
@@
  • If $g_{0}$ is quaternionic hyperbolic, then  [^3] Again, it is a statement of the theorem that the formula below is defined globally and not just locally.
@@
  • If $g_{0}$ is octonionic hyperbolic, then
    (3.5) $Q(S)=L\big(\frac{7}{24}\nabla^{*}\nabla S+\frac{1}{6}S+\frac{1}{3}\operatorname{tr}(S)g_{0}-\frac{1}{6}R^{\circ}(S)\big)-\frac{1}{24}\mathrm{Sym}(R^{\circ}(\nabla^{2}S)).$
Remark 1.
Note that using (2.18), the Hessian in (3.1) can be rewritten using the operator $\Pi_{4}$. This is what we will do in Subsection 3.2.
Note that for the real hyperbolic case, $Q(S)$ identifies with a symmetric $2$-tensor as
$\pi_{4}^{*}Q(S)=\pi_{2}^{*}\big(\frac{1}{4}\nabla^{*}\nabla S+\frac{1}{2}(\mathrm{tr}_{g_{0}}S)g_{0}-\frac{1}{2}S\big).$
This explains why the proof of the solenoidal injectivity is easier for this case.
@@
Consider a smooth path of metrics $(g_{\epsilon})$ and write $S=\partial_{\epsilon}g_{\epsilon}|_{\epsilon=0}$. First, we can identify all sphere bundles $S_{g}M$ by the pullback by $\Psi_{g}:SM\to S_{g}M$ defined by $\Psi_{g}(x,v):=(x,v/|v|_{g})$. Next, for $\epsilon$ small enough, let $\Psi_{\epsilon}\in C^{\alpha}(SM,SM)$ be the normal $(g_{\epsilon},g_{0})$-Morse correspondance (see [13, Section 3.2] for a precise definition) which gives an orbit equivalence:
(3.6) $\Psi_{\epsilon}(\varphi_{t}(z))=\varphi^{\epsilon}_{\kappa_{\epsilon}(z,t)}(\Psi_{\epsilon}(z))$
where the index $\epsilon$ indicates that we consider the corresponding objects for the metric $g_{\epsilon}$ and $\kappa_{\epsilon}$ is a time rescaling. Following Flaminio, we define
$\bar{\gamma}:(-\epsilon_{0},\epsilon_{0})\times\mathbb{R}\mapsto\bar{\gamma}(\epsilon,t):=\pi(\Psi_{\epsilon}(\varphi_{t}v)),$
where $\pi:SM\to M$ is the projection. This means that $\bar{\gamma}(\epsilon,\cdot)$ is a $g_{\epsilon}$ (non normalized) geodesic. We define $Y=\tfrac{d\bar{\gamma}}{dt}$ and $\xi=\tfrac{d\bar{\gamma}}{d\epsilon}$. Define $U(t,\epsilon):=U_{g_{\epsilon}}(\Psi_{\epsilon}(\varphi_{t}v))$ where $U_{g_{\epsilon}}(v)$ is the second fundamental form of the stable horosphere for the metric $g_{\epsilon}$ at $\pi_{\epsilon}(v)\in M$, seen a $(1,1)$-tensor. Note that, as already noticed in [13, Page 582] and with the notations of (2.22), along a geodesic, one has $UY_{j}=-\lambda_{j}Y_{j}$. Our starting point is [13, Proposition A, $(C_{3})$, Lemma 3.3.11, Lemma 4.2.1] which gives [^4] We use Pesin’s entropy formula to simplify the topological entropy appearing in [13, Proposition A, $(C_{3})$] with the sum of negative Lyapunov exponents appearing in [13, Lemma 3.3.11]:
(3.7) $\frac{d^{2}}{d\epsilon^{2}}\Phi(g_{\epsilon})|_{\epsilon=0}=\mathrm{Var}_{\mu_{\mathrm{Liou}}}^{g_{0}}(\operatorname{tr}(\nabla_{\xi}U)).$
Now, the goal is to write more explicitly $\operatorname{tr}(\nabla_{\xi}U)$ as a function $V(S)$ of the infinitesimal variation $S=\partial_{\epsilon}|_{\epsilon=0}g_{\epsilon}$. We now consider parallel fields $Y_{1}=Y,Y_{2},\ldots,Y_{n}$ as in (2.22). The Riccati equation satisfied by $U_{g}$ translates into a differential equation satisfied on the diagonal entries $B_{j}:=\langle(\nabla_{\xi}U)Y_{j},Y_{j}\rangle$. This is the content of [13, Lemma 4.3.1] and we have [^5] We are using a different convention for the sign of $R$, hence the sign difference with [13, Lemma 4.3.1]. :
(3.8) $\mathcal{L}_{Y}B_{j}-2\lambda_{j}B_{j}=-S(Y,Y)\lambda^{2}_{j}-g_{0}\left(\frac{\partial R_{\epsilon}}{\partial\epsilon}|_{\epsilon=0}(Y,Y_{j})Y,Y_{j}\right).$
@@
(3.9) $\sum_{j=1}^{n}B_{j}=\operatorname{tr}(\nabla_{\xi}U)\approx\frac{1}{2}\left(\sum_{j=2}^{n}\lambda_{j}\right)S(Y,Y)+\frac{1}{2}\sum_{j=1}^{n}\mu_{j}g_{0}\left(\frac{\partial R_{\epsilon}}{\partial\epsilon}|_{\epsilon=0}(Y,Y_{j})Y,Y_{j}\right),$
where $\mu_{1}=1$ and $\mu_{j}=\lambda_{j}^{-1}$ for $j>1$. Here, $\approx$ denotes the fact that the two functions are cohomologous, i.e., their difference is of the form $Xu$ where $X$ is the geodesic vector field associated to $g_{0}$ and $u\in C^{\infty}(SM)$ and we noted that the terms $\mathcal{L}_{Y}B_{j}$ are co-boundaries. We note that (3.8) for $j=1$ gives that $g_{0}\left(\frac{\partial R_{\epsilon}}{\partial\epsilon}|_{\epsilon=0}(Y,Y)Y,Y\right)\approx 0$ (recall from (2.22) that $\lambda_{1}=0$) and we can thus add it in the right hand side. Moreover, we can compute the diagonal entry $B_{1}$:
@@
To understand the second term in the right hand side of (3.9), we need to compute the variation of the $(3,1)$ curvature tensor defined in (2.20) with respect to the metric. By [26, Theorem 1.174], we have:
$\frac{\partial R_{\epsilon}}{\partial\epsilon}|_{\epsilon=0}S(X,Y)Z=(\nabla_{Y}{\nabla^{\prime}_{g}}S)(X,Z)-(\nabla_{X}{\nabla^{\prime}_{g}}S)(Y,Z)$
@@
$g_{0}({\nabla^{\prime}_{g}}S(X,Y),Z):=g_{0}\left(\frac{\nabla^{g_{\epsilon}}_{X}Y}{\partial\epsilon}|_{\epsilon=0},Z\right)=\frac{1}{2}(\nabla_{X}S(Y,Z)+\nabla_{Y}S(X,Z)-\nabla_{Z}S(X,Y)),$
where $S=\partial_{\epsilon}|_{\epsilon=0}g_{\epsilon}.$ For the sake of notations, we will denote by $C_{g_{0}}$ the $(3,0)$ tensor field
@@
$\sum_{j=2}^{n}\mu_{j}g_{0}\left(\frac{\partial R_{\epsilon}}{\partial\epsilon}|_{\epsilon=0}(Y,Y_{j})Y,Y_{j}\right)=\sum_{j=2}^{n}\mu_{j}\left(\nabla_{Y_{j}}(C_{g_{0}}S)(Y,Y,Y_{j})-\nabla_{Y}(C_{g_{0}}S)(Y_{j},Y,Y_{j})\right).$
@@
$\sum_{j=1}^{n}\mu_{j}g_{0}\left(\frac{\partial R_{\epsilon}}{\partial\epsilon}|_{\epsilon=0}(Y,Y_{j})Y,Y_{j}\right)\approx\sum_{j=1}^{n}\mu_{j}\nabla_{Y_{j}}(C_{g_{0}}S)(Y,Y,Y_{j}).$
@@
$\nabla_{Y_{j}}(C_{g_{0}}S)(Y,Y,Y_{j})=\nabla_{Y_{j},Y}^{2}S(Y,Y_{j})-\frac{1}{2}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y),$
where for a symmetric tensor $S$ and vector fields $X_{1},X_{2}$, one has
(3.10) $\nabla^{2}_{X_{1},X_{2}}S:=\nabla_{X_{1}}(\nabla_{X_{2}}S)-\nabla_{(\nabla_{X_{1}}X_{2})}S.$
We use the Ricci identity [26, Equation $(1.21)$] and (2.22), to re-write the first term:
@@
$=\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})-\lambda_{j}^{2}S(Y,Y).$
@@
$\frac{1}{4}\sum_{j=1}^{n}\mu_{j}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)=-\frac{1}{4}\sum_{j=1}^{n}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)=\frac{1}{4}\pi_{2}^{*}(\nabla^{*}\nabla S),$
@@
$\frac{1}{2}\sum_{j=1}^{n}\mu_{j}\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})=-\frac{1}{2}\sum_{j=1}^{n}\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})=\frac{1}{2}\pi_{2}^{*}(D_{g_{0}}D_{g_{0}}^{*}S)\approx 0$
@@
Putting everything together gives $V(S)=\pi_{4}^{*}Q(S)$ where $Q$ is defined in (3.2). Note that $Q$ is $4$-tensor which identifies, using (2.3), to the symmetric $2$-tensor $T$ used by Flaminio in [13, Proof of Proposition 1.3.3].
@@
$-\frac{1}{4}\sum_{j=1}^{n}\mu_{j}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)$ $=-\frac{1}{4}\sum_{j=1}^{n}\nabla_{Y_{j},Y_{j}}^{2}S(Y,Y)+\frac{1}{8}\nabla^{2}_{JY,JY}S(Y,Y)$
$=\frac{1}{4}\pi_{2}^{*}(\nabla^{*}\nabla S)+\frac{1}{8}\nabla^{2}_{JY,JY}S(Y,Y),$
@@
$\nabla^{2}_{JY,JY}S(Y,Y)=\pi_{4}^{*}\big((\nabla\nabla(S\circ J))\circ J\big)=\pi_{4}^{*}\big(\mathrm{Sym}((\nabla\nabla(S\circ J))\circ J)\big).$
@@
Here $\hat{X}_{i}$ denotes the $3$-tuple where we left out $X_{i}$. In particular one has, using (2.10) and (2.26),
@@
Now, using the fact that $J$ is parallel, we see that $\nabla_{v,Jv}S(v,Jv)=Y(\nabla S(Jv,v,Jv))$ is a co-boundary and we can thus ignore this term when computing the Hessian. Finally, we use (3.11) to obtain (3.3).
@@
by (2.28). This also shows that is term is independent of the choice of the base $(J_{1},J_{2},J_{3})$ and thus globally defined. The last term $\frac{1}{2}\sum_{j=1}^{n}\mu_{j}\nabla_{Y,Y_{j}}^{2}S(Y,Y_{j})$ is still a co-boundary by the exact same computation, using the fact that the $Y_{i}$ are parallel. We are left with understanding the middle term:
@@
$\mathrm{Sym}\left(\sum_{j=1}^{3}\nabla_{J_{j+1}Y,J_{j+1}Y}S(Y,Y)\right)=-\pi_{4}^{*}\left(\sum_{i=1}^{3}D_{g_{0}}^{J_{i}}D_{g_{0}}^{J_{i}}S\right).$
@@
Quotient of $\mathbb{O}\mathbb{H}^{n}$. This time the “trace” term in (3.11) gives
@@
where we used (2.53). The second to last term is still a coboundary and the remaining one is computed using using the fact that the $Y_{j}$ are parallel